OpenStudy (anonymous):

Find the value of the variable if P is between J & K. JP = 3y+1 ; PK = 12y - 4 ; JK = 75 Choose one answer. a. 5.2 b. 8.7 c. 3.9 d. 4.3

OpenStudy (anonymous):

Welcome to OpenStudy! Message me if you have any questions!

OpenStudy (anonymous):



OpenStudy (anonymous):

based on the above picture we get the following \[(3y+1) + (12y-4) = 75\] combine like terms and we get \[15y - 3 = 75\] add a 3 to both sides, we get \[15y = 78\] divide by 15 on both sides we have \[y = 78/15 = 5.2\]